A bowl contains 12 different flavored candies (4 chocolate, 5 strawberry and 3 menthol). A child takes 6 candies at random. What is the probability that the child took 2 of each flavor?


1
Expert's answer
2021-05-19T06:14:42-0400

Number of ways 6 candies to be chosen from 12 ="12C6"

The probability that the child took 2 of each flavor

"\frac{4C2*5C2*3C2}{12C6}"


"\frac{\frac{4*3}{2*1}*\frac{5*4}{2*1}*\frac{3*2}{2*1}}{\frac{12*11*10*9*8*7}{6*5*4*3*2*1}}"


"\frac{6*10*3}{11*4*3*7}"


"\frac{15}{77}"